# Bayesian and Multi-Objective Decision Support for Real-Time Cyber-Physical Incident Mitigation

### v1

@@ Abstract
This research proposes a real-time, adaptive decision-support framework for mitigating cyber incidents in cyber-physical systems, developed in response to an increasing reliance on these systems within critical infrastructure and evolving adversarial tactics. Existing decision-support systems often fall short in accounting for multi-agent, multi-path attacks and trade-offs between safety and operational continuity. To address this, our framework integrates hierarchical system modelling with Bayesian probabilistic reasoning, constructing Bayesian Network Graphs from system architecture and vulnerability data. Models are encoded using a Domain Specific Language to enhance computational efficiency and support dynamic updates. In our approach, we use a hybrid exposure probability estimation framework, which combines Exploit Prediction Scoring System and Common Vulnerability Scoring System scores via Bayesian confidence calibration to handle epistemic uncertainty caused by incomplete or heterogeneous vulnerability metadata. Mitigation recommendations are generated as countermeasure portfolios, refined using multi-objective optimisation to identify Pareto-optimal strategies balancing attack likelihood, impact severity, and system availability. To accommodate time- and resource-constrained incident response, frequency-based heuristics are applied to prioritise countermeasures across the optimised portfolios. The framework was evaluated through three representative cyber-physical attack scenarios, demonstrating its versatility in handling complex adversarial behaviours under real-time response constraints. The results affirm its utility in operational contexts and highlight the robustness of our proposed approach across diverse threat environments.
@@
cyber-physical systems , incident response , decision-support systems , multi-objective optimisation , Bayesian networks
journal: Computers & Security
@@ 1 Introduction
Cyber-Physical Systems (CPS) have become integral to daily life and infrastructure, enabling seamless interaction between digital and physical domains. Traditionally associated with Industrial Control Systems (ICSs), CPS are now extensively deployed in Connected and Autonomous Vehicles (CAVs), railway signalling infrastructures, and renewable energy networks. This expansion has drawn increased scrutiny toward cyber-physical security risks targeting critical infrastructure, from both attackers and defenders. In 2025, security researchers disclosed multiple vulnerabilities affecting solar photovoltaic (PV) inverters [10], highlighting how CPS components could be exploited to induce service outages, cause infrastructure damage, or in severe cases, inflict physical harm.
Given that many CPS deliver critical functions, real-time and adaptive decision support is essential for mitigating cyber incidents and minimising their impact. Here, “real-time” refers to the ability to assess and respond to threats within operational timeframes that preserve system safety and continuity, often under strict latency constraints. “Adaptive” refers to the capacity to dynamically adjust mitigation strategies based on evolving threat conditions, system states, and resource availability. However, several challenges remain. Unlike IT environments, where remediation typically involves deploying software patches or rebooting affected systems, CPS decisions must account for the interplay between cybersecurity and physical constraints [37] and take safety impacts into consideration.
In legacy systems and safety-critical applications, patching a vulnerable asset introduces significant operational risks, as poorly tested patches may lead to system failures or degraded functionality due to hardware limitations. Decision-makers must also determine the scope and extent of intervention: how much to mitigate, which assets to secure, and where to prioritise responses. These decisions are further complicated by the presence of multiple attack paths and adversary tactics, necessitating an integrated approach to countermeasure selection that balances disruption mitigation with system availability.
Recent decision-support systems (DSS) for CPS have introduced model-driven approaches leveraging simulation, system dynamics, and dependency modelling to guide risk mitigation planning [44]. While these systems support strategic assessment, they typically operate under static conditions and offer limited adaptability to evolving threats and system behaviours. For instance, most lack mechanisms to dynamically update model attributes, such as system properties or vulnerability exposure probabilities, during an ongoing cyber-physical incident. A typical example is the detection of a compromised communications dongle mid-response in a power grid, where the inability to update model parameters in real time reduces operational effectiveness in time-critical scenarios.
To address these limitations, our research adopts probabilistic methodologies to enable real-time CPS decision support. We hypothesise that effective CPS incident response strategy must be guided by data-driven risk metrics and time constraints. To test this, our framework builds on: Bayesian Network (BN) models for dynamic risk management [33]; threat modelling based on Failure Modes Effects and Criticality Analysis (FMECA) [3]; a hybrid exposure probability estimation framework that synergistically integrates Common Vulnerability Scoring System (CVSS) [30] and Exploit Prediction Scoring System (EPSS) [21] scores through a novel Bayesian confidence calibration technique; the use of Domain-Specific Languages (DSLs) for semantic interoperability and computational efficiency [17]; multi-objective optimisation, coupled with frequency-based heuristic analysis, to generate risk-informed countermeasure recommendations for decision-making [45]; and finally, adaptive refinement of the model, informed by iterative assessments of observed countermeasure effectiveness.
Figure 1: Architecture of our proposed framework
Our proposed framework, as depicted in Figure 1, is evaluated across three representative CPS scenarios: the 2015 Ukrainian power grid attack (BlackEnergy), a solar PV inverter system, and a railway Communication-Based Train Control (CBTC) system. Threat modelling for these scenarios was informed by the first author’s extensive cybersecurity experience in CPS, comprising 25 years of professional practice including 15 years as a Chief Information Security Officer. This depth of expertise ensured the inclusion of accurate technical insights and domain-relevant knowledge essential for credible CPS threat representation.
Contributions. Our main contributions are:
  • We propose a new way to leverage Bayesian Networks for modelling cyber attack scenarios in CPS environments. Unlike previous methods, our models are enriched with system-specific and vulnerability metadata, enabling probabilistic reasoning and risk-informed recommendations for real-time decision support. To ensure structured representation and computational efficiency, the models are encoded in a DSL format, which also facilitates dynamic updates to support adaptive decision-making in response to evolving threats.
@@
  • We use a hybrid exposure probability estimation framework that integrates the EPSS and the CVSS via a Bayesian confidence calibration technique. This framework yields robust and realistic recommendations under epistemic uncertainty stemming from incomplete or heterogeneous vulnerability metadata.
  • We propose a decision-support methodology for mitigating cyber incidents in CPS, centred on countermeasure portfolios derived through multi-
    objective optimisation and frequency-based heuristic analysis. As a novel contribution, mitigation probabilities are treated as parameters within a structured optimisation search space, enabling the generation of Pareto-optimal portfolios that form the foundation of the decision-support mechanism.
    frequency-based heuristic analysis across multiple optimisation runs is further employed to prioritise mitigation strategies under time and resource constraints. The methodology is evaluated across three representative CPS attack scenarios.
@@
  • The models and source code developed in this research, along with reference data and formulae used in the framework, are publicly available via the project’s GitHub repository [18]. This supports transparency, reproducibility, and further development by both research and practitioner communities.
@@
The remainder of this paper is structured as follows. Section 2 provides background on the key techniques adopted in this work. Section 3 outlines the proposed methodology, including its theoretical foundations and implementation workflow, illustrated with a running example. Section 4 presents further applications of the framework to CPS attack scenarios, accompanied by an analysis of experimental results. Section 5 provides an in-depth, technical discussion of the findings and outlines directions for future research. Section 6 reviews related research and identifies limitations in existing approaches. Finally, Section 7 summarises the key contributions of the study.
@@ 2 Background
Each of these nodes is associated with an exposure attribute which quantifies the likelihood that a CPS element is subject to exploitability conditions, failure dependencies, or cascading risk effects. As a critical factor in cybersecurity modelling, exposure shapes the propagation of vulnerabilities and attack vectors across interconnected CPS elements, influencing both the probability of attack success and the severity of resulting impacts. The exposure probability of each node is determined based on its attributes and its conditional dependence on parent nodes in the BN graph.
@@
CVSS is a standardised framework for assessing the technical severity of software vulnerabilities, with version 3.1 defining metrics to quantify exploitability and impact characteristics [15, 30]. However, it is a deterministic scoring system that does not account for the inherent uncertainty of real-world attacks. EPSS addresses this limitation by providing a probabilistic estimate of exploitation likelihood, grounded in empirical indicators such as exploit availability, vulnerability age, and active exploitation trends in the wild [21]. EPSS complements CVSS by introducing a dynamic, data-driven layer of exposure modelling that reflects temporal threat behaviour and captures uncertainty in real-world attacks.
@@
CPS owners and administrators are often cons-trained by fixed budgets that fall short of the minimum cost required for comprehensive system hardening. The challenge lies in selecting a subset of security measures that remain within budget while minimising residual risk from unaddressed vulnerabilities. Multi-objective optimisation addresses this challenge by identifying solutions that balance competing objectives, such as cost, system availability, and risk reduction. Prior research has explored Bayesian optimisation and evolutionary strategies to derive optimal security configurations under uncertainty and resource constraints [12, 29]. In particular, the work by Zebrowski et al. [45] formulates the optimisation task as a Pareto-efficient portfolio selection problem, leveraging a Bayesian Network to represent system vulnerabilities, mitigation actions, and operational constraints that define the feasible solution space. While this approach is highly applicable to CPS environments, it can be extended to incorporate mitigation strategies that account for cascading effects, adapt to evolving attacker tactics, and support time-sensitive decision-making in dynamic threat landscapes.
@@ 3 Methodology
This section describes the methodology of our proposed framework, illustrated with a running example of a solar PV inverter attack, which is based on the security reports by Dabrowski et al. [9] and Dashevskyi et al. [10]. In particular, we refer to their identified CVEs that can be chained to form hypothetical attack paths targeting solar PV systems. While Dabrowski et al. originally proposed destabilising the power grid via coordinated load modulation from a botnet of consumer devices, our scenario explores a complementary threat vector focused on the supply side. Here, the adversary leverages mobile or web applications to obtain sensitive information, which is then used to compromise a message broker and subsequently a controller linked to both the message broker and a solar PV inverter. Once initial access is established, the attacker hijacks additional inverters and synchronises their configuration changes to create abrupt fluctuations in grid input. By overwhelming the grid’s balancing mechanisms, this coordinated attack can degrade grid stability or induce power outages, thereby achieving the attacker’s ultimate objective.
Overview. Our framework (Figure 1) begins with the construction of a BN model that defines a shared structure for attack and impact graphs, informed by CPS system attributes and the attack scenario. Randomised mitigation probability values are introduced as initial inputs into a multi-objective optimisation process, which generates candidate countermeasure portfolios aimed at maximising system availability while minimising the likelihood of attack success and severe impact occurrence. These portfolios are subsequently reintegrated into the BN as updated mitigation probabilities, and the optimisation procedure is repeated across multiple trials within a single run. The top-performing portfolios from these iterative runs are then analysed using frequency-based heuristics to identify vulnerability nodes within the attack path that consistently exhibit high mitigation effectiveness. This insight serves to guide final prioritisation decisions during incident response, particularly when confronted with stringent time and resource limitations.
The high-level steps of our proposed framework are outlined in Algorithm 1. Each step is elaborated in detail in Sections 3.1 to 3.6.
Algorithm 1 Methodology of Proposed CPS Decision Support Framework
@@ 3.1 Bayesian Network Model Construction
In the construction of a representative and abstract BN model, nodes denote CPS assets, vulnerabilities, and hazards, and directed edges encode causal and conditional dependencies. The BN model is typically constructed by domain experts or system architects with deep knowledge of the CPS architecture and operational context. This process involves identifying key system components, mapping interdependencies, and integrating vulnerability data, often sourced from CVE repositories or expert elicitation [33].
This BN model serves as a shared structural foundation for both the Bayesian Attack Graph (BAG) and Bayesian Impact Graph (BIG) models. Leveraging the defined structure, the models integrate known system information alongside hypothesised attack scenarios to instantiate asset, vulnerability, and hazard nodes within the Bayesian framework, encoded in AutomationML [13].
@@
Figure 2 shows a Bayesian Network representation of a solar PV inverter attack scenario. The BN model is constructed by mapping system topology and functional interdependencies referenced from the work by Dashevskyi et al. [10]. CPS asset, vulnerability, and hazard nodes are instantiated using information from CVE repositories alongside the attack scenario [9].
Figure 2: Bayesian Network representation of a hypothetical cyber-physical attack on solar PV inverters connected to a power grid. Nodes shaded in yellow denote CPS assets; blue nodes represent vulnerabilities; and purple nodes correspond to hazards. The grey-shaded node indicates attack feasibility, while the red-shaded node reflects the attacker’s objective, which is to cause a power outage.
3.2 Exposure Calculations
This subsection outlines our methodology used to compute probabilities of attack success, which we term as exposures, for vulnerability, asset, and hazard nodes within the BN representation.
@@ 3.2 Exposure Calculations
To begin, the approach for vulnerability nodes needs to account for both publicly disclosed vulnerabilities with CVE identifiers, and proprietary or undocumented vulnerabilities lacking formal CVE enumeration. To do this, we employ an EPSS-based formulation for CVE-linked vulnerabilities, and a calibrated CVSS-based fallback for those without CVE identifiers. On the former EPSS-based formulation, the probability of exposure for a CVE-linked vulnerability is given by:
$P(E)_{\text{vuln, EPSS}}=\text{EPSS}$ (1)
where:
  • $P(E)_{\text{vuln, EPSS}}$ denotes the probability that the vulnerability will be exploited.
  • EPSS is the EPSS score retrieved from the FIRST EPSS database or API [16].
For vulnerabilities without CVE identifiers, we derive CVSS proxy estimates using the following formulation [15]:
$P(E)_{\text{vuln, CVSS}}=AV\times AC\times PR\times UI$ (2)
where each metric is assigned the corresponding numerical weight provided by FIRST [15].
Given that non-CVE vulnerabilities exhibit higher epistemic uncertainty due to limitations in CVSS scoring [40], we use a Bayesian confidence calibration technique, inspired by Kuppers et al. [28] to produce more realistic and robust exposure probability estimates. In our proposed formulation, both EPSS and CVSS scores are linked to a shared latent variable $P(E)^{*}_{\text{vuln}}$, representing the true, but unobserved, probability of exposure. These are modelled as distributions:
$P(E)_{\text{vuln, EPSS}}$ $\sim\mathcal{N}\left(P(E)^{*}_{\text{vuln}},\ \sigma^{2}_{\text{EPSS}}\right)$ (3)
$P(E)_{\text{vuln, CVSS}}$ $\sim\mathcal{N}\left(P(E)^{*}_{\text{vuln}},\ \sigma^{2}_{\text{CVSS}}\right)$ (4)
with $\sigma^{2}_{\text{CVSS}}>\sigma^{2}_{\text{EPSS}}$, reflecting the greater epistemic uncertainty associated with CVSS scoring.
Assuming a Gaussian prior as in [25], where $P(E)^{*}_{\text{vuln}}\sim\mathcal{N}(\mu_{0},\sigma_{0}^{2})$, the posterior distribution over $P(E)^{*}_{\text{vuln}}$ can be computed analytically as:
$\mu_{\text{post}}$ $=\frac{\frac{\mu_{0}}{\sigma_{0}^{2}}+\frac{P(E)_{\text{vuln, EPSS}}}{\sigma^{2}_{\text{EPSS}}}+\frac{P(E)_{\text{vuln, CVSS}}}{\sigma^{2}_{\text{CVSS}}}}{\frac{1}{\sigma_{0}^{2}}+\frac{1}{\sigma^{2}_{\text{EPSS}}}+\frac{1}{\sigma^{2}_{\text{CVSS}}}}$ (5)
$\sigma^{2}_{\text{post}}$ $=\left(\frac{1}{\sigma_{0}^{2}}+\frac{1}{\sigma^{2}_{\text{EPSS}}}+\frac{1}{\sigma^{2}_{\text{CVSS}}}\right)^{-1}$ (6)
In practice, since EPSS is unavailable for non-CVE vulnerabilities, the EPSS-related terms are omitted from the posterior update, and inference proceeds using only the prior and CVSS-derived $P(E)_{\text{vuln, CVSS}}$. The Bayesian latent variable formulation is thus represented as:
@@
The resulting calibrated exposure probability estimate $P(E)^{*}_{\text{vuln}}\sim\mathcal{N}(\mu_{\text{post}},\sigma^{2}_{\text{post}})$ serves as a unified, uncertainty-aware input for downstream attack propagation and mitigation optimisation in the BN graphs.
In addition, we incorporate an attack feasibility (AF) modifier, inspired by Xie et al. [41], in the formulation to better reflect real-world conditions in subsequent probability calculations for vulnerability nodes. Specifically, the AF modifier adjusts probability estimations based on system security posture, adversary capability, and environmental constraints. Systems with up-to-date defenses and restricted access tend to reduce the likelihood of a successful attack, resulting in lower AF values, whereas an unpatched CPS facing a skilled adversary will have a higher probability of successful attack. The final probability of successful attack of the vulnerability node is thus:
@@
  • $\phi$ is the attack feasibility factor, representing context-specific factors such as system defence posture and adversary capability.
@@
This formulation is grounded in Markov-based reliability analysis [26],
wherein asset failures are modelled as a stochastic process, capturing the progressive deterioration of system components over time. By incorporating time-dependent degradation patterns, this approach provides a probabilistic estimation of failure likelihood, supporting predictive maintenance and risk-informed decision-making in CPS.
@@
Hazard nodes, on the other hand, represent hazardous events, failures, or disruptions influenced by attack propagation. As such, the occurrence of a hazard is influenced by the cyber-physical dependencies within the system. If the parent node of a hazard is an asset, then the failure of that asset increases the likelihood of hazard activation. Similarly, if the parent node is another hazard, its occurrence propagates risk, potentially escalating the attack scenario.
@@
This binary assignment ensures deterministic modeling of attack escalation and structured attack progression through hazard nodes, where asset failures or hazard dependencies directly influence security outcomes.

### v2

# Bayesian and Multi-Objective Decision Support for Real-Time Incident Mitigation in Critical Infrastructure

@@ Abstract
Critical infrastructure increasingly relies on interconnected cyber-physical systems whose security incidents can escalate rapidly into safety and operational failures. Existing decision-support approaches struggle to support real-time incident response because they rely on static assumptions, incomplete vulnerability data, and single-objective risk models that do not adequately capture trade-offs between attack likelihood, impact severity, and system availability. This paper proposes a real-time, adaptive decision-support framework for incident mitigation in critical infrastructure that combines hierarchical system modelling with Bayesian probabilistic reasoning. The framework leverages probabilistic graphical models (Bayesian Networks) constructed from system architecture and vulnerability data, and employs confidence-calibrated exposure estimation to integrate complementary vulnerability scoring metrics under epistemic uncertainty. Mitigation strategies are explored as countermeasure portfolios and refined using multi-objective optimisation to identify Pareto-optimal trade-offs suitable for time- and resource-constrained response scenarios. Frequency-based heuristics are applied to prioritise robust mitigation actions across optimisation runs. The framework is evaluated on three representative cyber-physical attack scenarios, demonstrating its ability to adapt to evolving threats and provide actionable decision support under real-time constraints, thereby enhancing the operational resilience of critical infrastructure.
@@
critical infrastructure , cyber-physical systems , incident response , decision-support systems , multi-objective optimisation , Bayesian Networks
journal: Int. J. Crit. Infrastructure Prot.
@@ 1 Introduction
Cyber-Physical Systems (CPSs) form the backbone of modern critical infrastructure (CI), integrating digital and physical components to deliver the essential services that underpin modern society. While traditionally confined to centralised Industrial Control Systems (ICSs), CPSs are now deployed at scale across decentralised CI networks such as railway signalling networks and renewable energy ecosystems. This expansion has significantly broadened the attack surface for both state-sponsored and opportunistic actors, as digital vulnerabilities now translate directly into physical hazards and operational disruptions that threaten national security and public safety.
A pivotal example of this shifting threat landscape was highlighted in 2025 by the SUN:DOWN research report [11], which disclosed 46 new vulnerabilities across leading solar photovoltaic (PV) inverter vendors. These flaws, ranging from insecure cloud APIs to remote code execution, exposed severe systemic weaknesses in the renewable energy sector. Researchers demonstrated that by chaining these vulnerabilities, an adversary could orchestrate a coordinated load-changing attack, manipulating inverter output to destabilise grid frequency. Such an exploit could necessitate large-scale load shedding or trigger emergency equipment shutdowns, underscoring the existential risk that unsecured CPS components pose to infrastructural resilience.
Given the critical nature of these processes, real-time and adaptive decision support is essential to ensure robust security and incident mitigation, while simultaneously minimising impacts on safety and preserving continuous system availability. In this context, “real-time” refers to the decision-support window within which the CI operator needs to respond to preserve system stability, often under stringent latency constraints. “Adaptive” refers to the capacity to dynamically adjust mitigation strategies based on evolving threat conditions, system states, and resource availability. Unlike traditional IT environments, where remediation typically involves rapid software patching, CPS security decisions must balance cybersecurity objectives against physical process constraints [40]. In legacy CI, applying an unvalidated patch entails substantial operational risk. Updates may induce latent system failures or performance degradation, necessitating countermeasures that prioritise the resilience of CI systems [35].
Despite advancements in model-driven decision support, existing frameworks often operate under static assumptions and lack the adaptability required during an unfolding incident [29, 47]. Most frameworks cannot update model parameters such as vulnerability exposure probabilities or asset interdependencies in real time. A representative scenario is the detection of a compromised communication dongle mid-response in a power grid. The inability to dynamically revise model attributes constrains operational decision-making and reduces response efficacy in time-critical contexts.
To address these limitations, we propose a real-time, adaptive decision-support framework for incident mitigation in critical infrastructure, grounded in applied probabilistic reasoning and designed to meet the operational realities of CI protection. We posit that effective response strategies must be guided by data-driven risk metrics while remaining bounded by stringent time and resource constraints. As illustrated in Figure 1, the proposed framework integrates Bayesian Network (BN) models for dynamic risk management [34, 48] with threat modelling informed by Failure Modes, Effects, and Criticality Analysis (FMECA) [3]. The key novelty lies in the operationalisation of confidence-calibrated exposure metrics, which bridge the gap between theoretical risk modelling and the practical demands of real-world CI incident response.
Figure 1: Architecture of our proposed critical infrastructure security framework. The diagram shows an iterative process integrating a CPS model (vulnerabilities, hazards, assets, exposure/impact probability attributes) with BN construction (BAG and BIG). Multi-objective optimisation generates Pareto-optimal countermeasure portfolios, updating model attributes and propagating exposure/impact probabilities. Frequency heuristic analysis selects the optimal portfolio.
Contributions. The main contributions of our paper are summarised below:
  • We present a Bayesian Network modelling approach designed explicitly for real-time, adaptive incident response in CI, rather than offline risk analysis. The model embeds system and vulnerability metadata in a Domain-Specific Language (DSL) [17], enabling rapid updates to structure and parameters as incidents unfold.
@@
  • We develop a hybrid exposure probability estimation framework that integrates complementary vulnerability scoring metrics through a Bayesian confidence calibration technique. This provides robust recommendations even under epistemic uncertainty stemming from incomplete or heterogeneous vulnerability metadata.
  • We introduce a multi-objective decision-support methodology for incident mitigation in cyber-physical systems that balances security, safety, and operational resilience. By treating mitigation probabilities as dynamic decision variables, the framework identifies Pareto-optimal countermeasure portfolios tailored to CI constraints, and uses frequency-based heuristics to prioritise robust mitigation strategies under time- and resource-critical conditions.
  • We demonstrate the applicability of the proposed framework through three representative CI case studies: the 2015 Ukrainian power grid (BlackEnergy) attack, a solar PV inverter system, and a railway CBTC network, demonstrating its effectiveness across diverse CI domains.
@@
  • The models and source code developed in this research, along with the reference data and formulae used in the framework, are publicly available via the project’s GitHub repository [20]. This supports transparency, facilitates reproducibility and enables further development by both the research and practitioner communities within the CI protection domain.
@@
The remainder of this paper is structured as follows. Section 2 provides background on the key techniques adopted in this work. Section 3 outlines the proposed methodology, including its theoretical foundations and implementation workflow, illustrated with a running example. Section 4 presents further applications of the framework to CI attack scenarios, accompanied by an analysis of experimental results. Section 5 provides an in-depth, technical discussion of the findings and outlines directions for future research. Section 6 reviews related research and identifies limitations in existing approaches. Finally, Section 7 summarises the key contributions of the study.
@@ 2 Background
Each node is associated with exposure and impact probability attributes that quantify the likelihoods of exploitation, failure dependencies, and cascading risks. These attributes shape the propagation of vulnerabilities and attack vectors, influencing the probabilities of attack success and of severe impact to the CI. Our work builds on these foundations by integrating such BN-based exposure and impact modelling into a real-time decision-support loop for incident mitigation, and by coupling the BN with optimisation and heuristic prioritisation to drive operational countermeasure selection rather than using the BN solely for static risk analysis.
@@
The Common Vulnerability Scoring System
(CVSS) provides a standardised framework for assessing the technical severity of vulnerabilities, with version 3.1 defining metrics to quantify exploitability and impact characteristics [16, 30]. However, as a deterministic scoring system, CVSS often fails to account for the inherent uncertainty and temporal dynamics of real-world attacks. The Exploit Prediction Scoring System (EPSS) addresses this limitation by providing probabilistic estimates of exploitation likelihood over the subsequent 30-day window, grounded in empirical indicators such as exploit availability and active exploitation trends [21].
To achieve a more granular assessment, our framework further integrates the Known Exploited Vulnerabilities (KEV) catalogue [9], which identifies vulnerabilities confirmed to have been exploited in real-world scenarios, alongside Likely Exploited Vulnerabilities (LEV) [31] data, which provides daily probability estimates of observed exploitation activities. By incorporating both authoritative records and statistical exploitation estimates, our approach captures a broader spectrum of the threat landscape.
Our framework distinguishes itself by fusing these complementary scoring schemes within a Bayesian calibration mechanism, yielding uncertainty-aware exposure probabilities that explicitly account for incomplete or heterogeneous vulnerability metadata, and by embedding these calibrated probabilities directly into the BN that underpins our decision-support process.
@@
CI decision makers are often constrained by fixed budgets that fall short of the minimum cost required for comprehensive system hardening. The challenge lies in selecting a subset of security measures that remain within budget while minimising residual risk from unaddressed vulnerabilities [13]. Multi-objective optimisation addresses this challenge by identifying solutions that balance competing objectives, such as cost, system availability, and risk reduction. Prior research has explored Bayesian optimisation and evolutionary strategies to derive optimal security configurations under uncertainty and resource constraints [13, 29]. In particular, the work by Zebrowski et al. [48] formulates the optimisation task as a Pareto-efficient portfolio selection problem, leveraging a Bayesian Network to identify security measures that simultaneously minimise various types of expected cyberattack impacts, while satisfying budgetary and other constraints. While this approach is highly applicable to CI environments, our work extends it by treating mitigation probabilities as parameters that shape the decision variables within the optimisation space, repeatedly updating BN parameters in response to candidate portfolios, and applying frequency-based heuristic analysis across multiple optimisation runs to prioritise countermeasures under time- and resource-constrained, real-time incident response conditions.
@@ 3 Methodology
This section describes the methodology of our proposed framework, illustrated with a running example of a solar PV inverter attack, which is based on the security reports by Dabrowski et al. [10] and Dashevskyi et al. [11]. In particular, we refer to their identified CVE (Common Vulnerabilities and Exposures) entries that can be chained to form hypothetical attack paths targeting solar PV inverter systems. While Dabrowski et al. originally proposed destabilising the power grid via coordinated load modulation from a botnet of consumer devices, our scenario explores a complementary threat vector focused on the supply side. Here, the adversary leverages mobile or web applications to obtain sensitive information, which is then used to compromise a message broker and subsequently a controller linked to both the message broker and a solar PV inverter. Once initial access is established, the attacker hijacks additional inverters and synchronises their configuration changes to create abrupt fluctuations in grid input. By overwhelming the grid’s balancing mechanisms, this coordinated attack can degrade grid stability or induce power outages, thereby achieving the attacker’s ultimate objective.
Overview. Our framework (Figure 1) begins with the construction of a BN model that defines a shared structure for attack and impact graphs, informed by CPS system attributes and the attack scenario. Randomised mitigation probability values are introduced as initial inputs into a multi-objective optimisation process, which generates candidate countermeasure portfolios aimed at maximising system availability while minimising the likelihood of attack success and severe impact occurrence. These portfolios are subsequently reintegrated into the BN as updated mitigation probabilities, and the optimisation cycle is repeated across multiple trials within a single run. The top-performing portfolios from these iterative runs are then analysed using frequency-based heuristics to identify vulnerability nodes within the attack path that consistently exhibit high mitigation effectiveness. This insight serves to guide final prioritisation decisions during incident response, particularly when confronted with stringent time and resource limitations. The high-level steps of our proposed framework are outlined in Algorithm 1. Each step is elaborated in detail in Sections 3.1 to 3.6.
Algorithm 1 Methodology of Proposed CI Decision Support Framework
@@ 3.1 Bayesian Network Model Construction
In the first step of our methodology, we construct an abstract BN model that captures the relevant cyber-physical structure of the target CPS and the incident under consideration. Nodes represent CPS assets, vulnerabilities, or hazards, while directed edges encode the causal and conditional dependencies through which attacks and failures may propagate. The model is instantiated using system-specific architectural information and operational knowledge, typically elicited from domain experts or system documentation, together with vulnerability data drawn from CVE repositories or expert judgement.
This BN serves as a shared structural backbone for both the Bayesian Attack Graph (BAG) and the Bayesian Impact Graph (BIG) used in subsequent stages of the framework. Given the fixed structure, we incorporate scenario-specific assumptions and hypothesised attack paths to populate asset, vulnerability, and hazard nodes, and encode the resulting model in AutomationML (XML based data modeling language for representing engineering models) [14] to support systematic analysis and downstream optimisation.
@@
Figure 2 shows a Bayesian Network graph of a solar PV inverter attack scenario. The BN model was constructed by mapping system topology and functional interdependencies referenced from the work by Dashevskyi et al. [11]. CPS asset, vulnerability, and hazard nodes are instantiated using information from CVE repositories alongside the attack scenario [10].
Figure 2: Bayesian Network of a hypothetical cyber-physical attack on solar PV inverters connected to a power grid. Node colours denote assets (orange), vulnerabilities (green), hazards (yellow), attack feasibility (cyan), and attacker goal (red). A high-resolution version of the graph is available in the project’s GitHub repository [20].
3.2 Exposure Probability Calculations
This subsection outlines our methodology used to compute probabilities of attack success, which we term as exposure, for vulnerability, asset, and hazard nodes within the BN representation.
@@ 3.2 Exposure Probability Calculations
The approach for vulnerability nodes needs to account for both publicly disclosed vulnerabilities with CVE identifiers, and proprietary or undocumented vulnerabilities lacking formal CVE enumeration. To calculate the exposures of CVE-linked vulnerability nodes, we use EPSS, LEV, and KEV metrics in conjunction with the maximal-envelope approach introduced by Mell et al. [31] to select the highest exposure probability among these complementary vulnerability scoring frameworks.
The Likely Exploited Vulnerability (LEV) metric is defined by Mell et al. as:
$\text{LEV}(v,d_{0},d_{n})>=1-\prod_{\forall d_{i}\in\text{dates}(d_{0},d_{n},30)}\left(1-\text{epss}(v,d_{i})\times\text{weight}(d_{i},d_{n},30)\right)$ (1)
where $v$ is the vulnerability; $d_{0}$ is the first date on which an EPSS score is available; $d_{n}$ is the date on which the calculation is performed (usually the present day); $\text{epss}(v,d)$ is the EPSS score for vulnerability $v$ on date $d$; and $\text{weight}(d_{i},d_{n},w)$ is the number of days between $d_{i}$ and $d_{n}$, limited by the window size $w$, divided by $w$.
The probability of exposure for a CVE-linked vulnerability, $P(E)_{\text{vuln, CVE}}$ is defined as:
$P(E)_{\text{vuln, CVE}}=\max\left(\text{EPSS, LEV, KEV}\right)$ (2)
where $\text{KEV}\in\{0,1\}$ acts as a logical ceiling, ensuring that any vulnerability confirmed to have been exploited in real-world scenarios is assigned a maximal exposure probability of 1; otherwise, it is assigned 0. Notably, if $KEV=1$, $P(E)_{\text{vuln,CVE}}$ becomes 1 regardless of EPSS or LEV, as this is a critical safety-conservative assumption for CI.
For proprietary or undocumented vulnerabilities lacking CVE identifiers, EPSS, LEV, and KEV metrics are unavailable for exposure probability calculation. Proxy CVSS scores are therefore derived from the exploitability subscore of the CVSS base metrics, computed as follows:
$P(E)_{\text{vuln, non-CVE}}=AV\times AC\times PR\times UI$ (3)
where $AV$, ACAC, PRPR, and $UI$ represent the Attack Vector, Attack Complexity, Privileges Required, and User Interaction metrics, respectively. Each metric is assigned the corresponding numerical weight per the CVSS v3.1 specification [16].
As shown by Wunder et al. [43], epistemic uncertainty persists in CVSS evaluations [43]. To address this, we employ a Bayesian confidence calibration technique inspired by Küppers et al. [28]. We model the observed proxy CVSS score as a distribution linked to a shared latent variable $P(E)^{*}_{\text{vuln}}$, representing the unobserved true probability of exposure:
$P(E)_{\text{vuln, non-CVE}}$ $\sim\mathcal{N}\left(P(E)^{*}_{\text{vuln}},\sigma^{2}_{\text{CVSS}}\right)$ (4)
Assuming a Gaussian prior as in [25], where $P(E)^{*}_{\text{vuln}}\sim\mathcal{N}(\mu_{0},\sigma_{0}^{2})$, the Bayesian latent variable formulation for the posterior parameters is represented as:
@@
The resulting calibrated exposure probability estimate $P(E)^{*}_{\text{vuln}}\sim\mathcal{N}(\mu_{\text{post}},\sigma^{2}_{\text{post}})$, serves as a unified, uncertainty-aware input for downstream attack propagation and mitigation optimisation within the BN graphs.
In addition, we incorporate an Attack Feasibility (AF) modifier, inspired by the work of Xie et al. [44], to ensure the formulation reflects real-world conditions more accurately. This modifier is applied to the probability calculations for all vulnerability nodes. Specifically, the AF modifier adjusts probability estimations by accounting for the system’s security posture, adversary capabilities, and environmental constraints. For instance, systems featuring robust, up-to-date defences and restricted access protocols yield lower AF values, reflecting a reduced likelihood of exploitation. Conversely, an unpatched CPS targeted by a sophisticated adversary results in a higher probability of a successful attack. Consequently, the final probability of a successful attack for a given vulnerability node, irrespective of whether it has a CVE identifier, is defined as:
@@
  • $\phi$ is the attack feasibility (AF) modifier, representing context-specific variables such as system defence posture, adversary capability, and environmental constraints.
@@
This formulation is grounded in Markov-based reliability analysis [26], in which asset failures are modelled as a stochastic process, capturing the progressive deterioration of system components over time. By incorporating time-dependent degradation patterns, this approach provides a probabilistic estimation of failure likelihood, supporting predictive maintenance and risk-informed decision-making in CI environments.
@@
Finally, hazard nodes represent hazardous events, failures, or disruptions influenced by attack propagation. As such, the occurrence of a hazard is influenced by the cyber-physical dependencies within the system. If the parent node of a hazard is an asset, then the failure of that asset increases the likelihood of hazard activation. Similarly, if the parent node is another hazard, its occurrence propagates risk, potentially escalating the attack scenario.
@@
This binary assignment ensures deterministic modelling of attack escalation and structured attack progression through hazard nodes, where asset failures or hazard dependencies directly influence security outcomes.